Two charges 9e and 3e are placed at a certain distance r.The distance of the point where the electric field intensity will be zero is:
 

Grade:12

1 Answers

Soumya Ranjan Mohanty
6 years ago
Dear student,
Let distance at which electric field intesity is zero be 'a' from 9e charge.
so putting formula, the electric field for 3e and 9e will be equal,
1/(4*pi*epsilon0) *(9e/a^2)=1/(4*pi*epsilon0)* (3e/(r-a)^2)
=>3(r-a)^2 = a^2
solving we get,
a= r*sqrt(3)/(sqrt(3)+1)= r*(3-sqrt(3))/2 [Ans]
Regards,
Thank you,
Soumya Ranjan Mohanty.